OpenStudy (anonymous):

Use transformations and the zeros of the polynomial function f(x) = x(x+3)(x-6) to determine the zeros of each of the following functions.

OpenStudy (anonymous):

c) y = f(1/2 x) and d) y=f(2x)

OpenStudy (jlastino):

Just replace the x in the factors w/ 1/2 and 2x respectively and equate each to 0 you'll find the zeroes

OpenStudy (anonymous):

ohh, i see. alright, thank you!

OpenStudy (jlastino):

c. (1/2)x +3 = 0 Solve for x (1/2)x -6 = 0 Solve for x d. 2x+3=0 Solve for x 2x-6=0 Solve for x

OpenStudy (anonymous):

by why is the answer for c) 0, -6, 12? D:

OpenStudy (jlastino):

there should only be 2 values of x.... 1/2x =-3 x=-6 1/2x = 6 x=12

OpenStudy (anonymous):

but the answer from the back of my text also has 0..

OpenStudy (jlastino):

oh i forgot there's an x haha right

OpenStudy (jlastino):

so 0 (0+3)(0-6) =0

OpenStudy (anonymous):

what do you mean? o.o

OpenStudy (jlastino):

the given equation is x(x+3)(x-6) right? you are going to look for values of x that would make the resulting value of those =0

OpenStudy (anonymous):

yes. ohh, yup.

OpenStudy (jlastino):

so for letter c, y(1/2x) (1/2)x)((1/2x)+3)((1/2)x-6) when we make x=0 1/2(0) Judging by the first factor ....we can already say that the resulting value is 0 because anything multiplied to zero would be zero :D

OpenStudy (anonymous):

ohh, okay. thank you!